OpenStudy (anonymous):

3x + y = 6 6x + 2y = 8 ????

OpenStudy (anonymous):

6x + 2y = 8 3x + y = 6 (2) 6x + 2y = 8 subtract 6x + y = 12 y= -4

OpenStudy (anonymous):

now substitute -4 for y and figure out x